You find the following corporate bond quotes. To calculate the number of years until maturity, assume that it is currently Ja What is the yield to maturity for the bond issued by Xenon, Inc.?
Answer #1

Par Value =2000
Coupon =2000*7.30% =146
Price =94.411/100*2000 =1888.22
Number of years =25
YTM using excel formula=Rate(25,146,-1888.22,2000) =7.82%